# Aspects of the chiral crossover transition in (2+1)-flavor QCD with Möbius domain-wall fermions

### v1

@@ IV Results
IV.1 Estimating the pseudo-critical temperature
@@
for which we present our results in Fig. 1. Since the transition associated with the restoration of the non-singlet $\mathrm{SU}_{A}(2)$ symmetry is a smooth crossover in QCD with physical quark masses, this observable has a rather smooth dependence on the temperature. The chiral susceptibility, on the other hand, is known to be more sensitive to the restoration of symmetries and, in particular, to diverge in the chiral limit. This also holds true for its disconnected part, whereas its connected part is expected to diverge only if the $\mathrm{U}_{A}(1)$ symmetry is restored.
FIG. 1: The subtracted chiral condensate $\Delta_{ls}$ normalized by $(T/T_{pc})^{4}$ is shown as a function of the temperature $T$ calculated using MDW fermions (red pentagons) on an $N_{\tau}=8$ lattice and compared with results from Ref. [55] obtained using the HISQ discretization on $N_{\tau}=8,12,16$ lattices. For the normalization factor $T_{pc}$ we used the pseudo-critical temperature $T_{pc}=155\,\mathrm{MeV}$ reported in the earlier work [9].
The results for the disconnected part of the chiral susceptibility as a function of the temperature $T$ is shown in Fig. 2 (top). This observable shows a pronounced peak in a small range of temperatures and falling off both at lower and higher temperatures. The location of this peak defines a pseudo-critical temperature, which we denote as $T_{pc,N_{\tau}=8}$.
For the determination of the pseudo-critical temperature, we normalized the data points of the disconnected part of the chiral susceptibility in the temperature-independent form $m_{s}^{2}\,\chi_{\mathrm{disc}}/T_{pc}^{4}$, using the temperature values determined using the $r_{0}$-scale and the pseudo-critical temperature $T_{pc}$ from Refs. [9, 52], which are also given in Table 1. The normalized data points are interpolated using a Padé fit of order $[1/3]$. From the zero of the derivative of this rational function with respect to the temperature, shown in Fig. 2 (bottom), we have estimated the pseudo-critical temperature to be
@@
which we henceforth refer to as $T_{pc}$. This is consistent with the earlier estimate of the pseudo-critical temperature, $155(9)\,\mathrm{MeV}$, given in Ref. [9], with the improvement of reduced statistical errors. Our estimate of $T_{pc}$ is also consistent with the corresponding ones obtained using the HISQ and stout actions after performing a continuum extrapolation, which are $156.5(1.5)\,\mathrm{MeV}$ [4] and $158.0(6)\,\mathrm{MeV}$ [6], respectively. Furthermore, $\chi_{\text{disc}}$ becomes small only at temperatures $T>186\,\mathrm{MeV}$, signalling a significant contribution from the topological fluctuations of the QCD vacuum in the vicinity of $T_{pc}$ which we discuss in detail in the next section.
FIG. 2: The disconnected part of the chiral susceptibility shown as a function of $T$ (top) and the extraction of the pseudo-critical temperature $T_{pc,N_{\tau}=8}$ (black point) from the point where the slope of $m_{s}^{2}\chi_{\text{disc}}$ goes to zero (bottom). For the normalization factor $T_{pc}$ we used the pseudo-critical temperature $T_{pc}=155\,\mathrm{MeV}$ reported in the earlier work [9].
@@
Results for this susceptibility are shown in Fig. 3. While the statistical noise on this observable is relatively large, it also shows a peak at $T_{N_{\tau}=8}^{\chi_{us}}\simeq 160\,\mathrm{MeV}$ within its quite large statistical errors, which is close to the pseudo-critical temperature extracted earlier using $\chi_{\mathrm{disc}}$, reaffirming the fact that even with physical quark masses the (2+1)-flavor QCD respects $\mathrm{SU}_{L}(2)\times\mathrm{SU}_{R}(2)$ symmetry to a very good extent. It was already noted earlier [56] that the strange quark mass is relatively heavy and does not break the almost exact two flavor chiral symmetry in the QCD action due to the light quarks explicitly. It therefore appears as an external parameter in the energy-like scaling variable used in the scaling analysis in the vicinity of the chiral transition, just like the chemical potentials corresponding to the conserved charges. The estimated $T_{\text{HISQ}}^{\chi_{us}}=159.39(22)$ [56] using (2+1)-flavor HISQ gauge ensembles on an $N_{\tau}=8$ lattice is consistent with our MDWF results.
@@
FIG. 3: The mixed-mass susceptibility $\chi_{us}$ shown as a function of the temperature $T$. For the MDWF results the position of the peak was determined at a temperature of $T_{N_{\tau}=8}^{\chi_{us}}=160\,\mathrm{MeV}$. We have normalized with factor $T_{pc}=158.7$ MeV given in Eq. (14).
@@
Having discussed the restoration of the non-singlet part of chiral symmetry, it would be also interesting to check how well the chiral Ward identities are followed by the MDWF action, even at finite lattice spacing.
IV.2 The fate of the singlet and non-singlet parts of chiral symmetry
@@
Once the non-singlet $\mathrm{SU}_{A}(2)$ part of chiral symmetry is restored at $T_{pc}$, one would then ask the question, when the singlet $\mathrm{U}_{A}(1)$ sub-group of the chiral symmetry is effectively restored. We now want to estimate the magnitude of the $\mathrm{U}_{A}(1)$ breaking at temperatures greater than $T_{pc}$, which is done through a calculation of the observable $\chi_{\pi}-\chi_{\delta}$ as motivated earlier. According to the chiral Ward identities, the integrated pion correlator $\chi_{\pi}$ should be related to the chiral condensate as seen in Eq. (7). On the other hand, the connected piece of the chiral susceptibility is just the scalar iso-triplet correlator, $\chi_{\text{conn}}\equiv\chi_{\delta}$. Previous studies [22, 45] with domain-wall fermions have reported $\chi_{\pi}-\chi_{\delta}$ to be finite at $196\,\mathrm{MeV}$, anticipating that the $\mathrm{U}_{A}(1)$ symmetry will be effectively restored beyond this temperature [9]. A different study, also with another variant of domain-wall fermions, but on relatively small lattice volumes, reports on the effective restoration of the $\mathrm{U}_{A}(1)$ symmetry at temperatures around $T_{c}$ [57].
@@
In order to address this apparently unsettled issue, we calculate the temperature dependence of the topological susceptibility to have an independent measure of the $\mathrm{U}_{A}(1)$ breaking. We recall that the restoration of the non-singlet part of chiral symmetry can also be inferred from Eq. (12). Moreover, the finite volume correction to $\chi_{\mathrm{top}}$ is dominated by the $\eta^{\prime}$ meson mass rather than the pion, and is expected to be mild at temperatures beyond $T_{pc}$. We obtained $\chi_{\text{top}}$ on the gauge configurations generated with MDWF, by calculating the average variance of the number of zero modes counted using as probe the overlap Dirac operator, details of which are mentioned in Section III. Our results for $\chi_{\text{top}}$ as a function of the temperature are shown in Fig. 4. From next-to-leading order (NLO) chiral perturbation theory, the topological susceptibility can be derived to be
$\chi_{\text{PT}}^{\text{NLO}}$ $=\frac{z}{(1+z)^{2}}M^{2}_{\pi}F^{2}_{\pi}[1+\delta_{1}],\;$ (16)
@@
where $\delta_{1}=\frac{2M_{\pi}}{F^{2}_{\pi}}(h_{1}^{r}-h_{3}^{r}-l_{4}^{r}-l_{7})$ in terms of low energy constants of SU(2) chiral perturbation theory [58]. For our case, $z$, which is the ratio of mass of up and down quarks, is simply $z=1$. Using the latest values of $F_{\pi}=94.15\,\mathrm{MeV}$, $M_{\pi}=134.98\,\mathrm{MeV}$ and the quantity $(h_{1}^{r}-h_{3}^{r}-l_{4}^{r}-l_{7})=-0.0114$ from the FLAG review [59], we obtain the $(\chi_{\text{PT}}^{\text{NLO}})^{1/4}=78.8\pm 2.9\,\mathrm{MeV}$, where the error predominantly comes from the error in the determination of $F_{\pi}$. As discussed in the previous section, the contribution to the chiral observables from strange quarks is small compared to the light quarks. Hence, we do not consider three-flavor chiral perturbation theory. Our results for the topological susceptibility in (2+1)-flavor QCD (triangles) with MDWF start to deviate from the NLO chiral perturbation theory prediction shown as a band in Fig. 4 below $T_{pc}$, alluding towards an effective restoration of the non-singlet part of chiral symmetry. The deviation from the chiral perturbation theory result becomes larger with increasing temperatures, showing a $\approx 50\%$ deviation at the highest temperature of $186\,\mathrm{MeV}$. However the magnitude of $\chi_{\text{top}}$ at temperatures higher than $T_{pc}$ successively gets closer to the values of $m_{\ell}^{2}\,\chi_{\text{disc}}$, besides a small deviation resulting from the finite mass of the light quarks used in this work, signalling an effective restoration of the non-singlet part of chiral symmetry.
@@
Furthermore, the temperature dependence of our data for $\chi_{\text{top}}^{1/4}/T_{pc}$ at $T>160\,\mathrm{MeV}$ can be characterized by $(T/T_{pc})^{-\alpha}$ with $\alpha=2.8(1)$. The exponent $\alpha\approx 2$ describes the temperature dependence of the topological susceptibility in QCD with two light quark flavors within a dilute instanton gas model, which assumes fluctuations of isolated uncorrelated instantons, and should persist at asymptotically high temperatures. The fact that the exponent $\alpha>2$, reflects a non-trivial breaking of the $\mathrm{U}_{A}(1)$ symmetry up to $T\approx 186\,\mathrm{MeV}$, which is unlikely due to a dilute instanton gas-like scenario. It is also known from a recent study with the HISQ discretization that the $\mathrm{U}_{A}(1)$ symmetry is effectively restored at $T\gtrsim 180\,\mathrm{MeV}$, after performing a continuum extrapolation [32]. This is consistent with the fact that we do not observe an effective restoration of the $\mathrm{U}_{A}(1)$ part of chiral symmetry up to the highest temperature we have studied using MDWF, which is $T=186\,\mathrm{MeV}$.
IV.3 Comparing Möbius domain-wall results with highly-improved staggered quark data
@@
Having obtained precise results on the chiral observables using MDWF, even on a finite $N_{\tau}=8$ lattice, it is imperative to determine how well these results compare with results obtained using different fermion discretizations. The staggered fermion discretization scheme explicitly breaks the non-singlet flavor symmetries on the lattice and only respects a $\mathrm{U}(1)$ sub-group of the whole chiral symmetry group of QCD with two light quark flavors, which is an admixture of the physical flavor and the taste symmetries. It is known that the chiral symmetries in terms of the chiral Ward identities are recovered in the continuum limit for observables calculated using the HISQ discretization. This has been shown, for example, through the exact matching of the disconnected part of the chiral susceptibility and the topological susceptibility in the continuum limit for physical gauge ensembles in (2+1)-flavor QCD generated with the HISQ action [27, 32]. On the other hand, in earlier work with MDWF in Ref. [9], it was observed that $\chi_{\text{disc}}$ calculated using domain-wall fermions was $50\%$ larger than the result computed using the staggered discretization in the crossover region. It would thus be important to now address this puzzle, especially since we know that for a meaningful comparison one has to perform a continuum extrapolation of the staggered results.
@@
FIG. 5: $m_{s}^{2}\,\chi_{\mathrm{disc}}/T_{pc}^{4}$ computed with the MDWF action for $N_{\tau}=8$, in comparison to results obtained using the HISQ action. The continuum estimate was computed from the HISQ $N_{\tau}=12,16$ results only. For the normalization factor $T_{pc}$ we used the pseudo-critical temperature $T_{pc}\equiv T_{pc,N_{\tau}=8}$ given in Eq. (14).
@@
Results for the observable $\chi_{\text{disc}}$ as a function of the temperature are shown in Fig. 5, comparing the results obtained using the HISQ discretization for lattice spacings corresponding to $N_{\tau}=8,12,16$ with our results calculated using MDWF on an $N_{\tau}=8$ lattice. From a continuum extrapolation of the peak positions associated with the HISQ $N_{\tau}=8,12,16$ data for $\chi_{\mathrm{disc}}$, we have estimated the pseudo-critical temperature with an $\mathcal{O}(1/N_{\tau}^{2})$ fit resulting in $T_{pc}^{\mathrm{HISQ}}=157.4(7)\,\mathrm{MeV}$, which agrees well within errors with our corresponding MDWF $N_{\tau}=8$ estimate of $T_{pc,N_{\tau}=8}=158.7{}_{{}\mathop{-}2.3}^{{}\mathop{+}2.6}\,\mathrm{MeV}$, as given in Eq. 14. Unlike the location of its peaks, the numerical values of $\chi_{\mathrm{disc}}$ obtained using HISQ action on an $N_{\tau}=8$ lattice are clearly not within the $\mathcal{O}(a^{2})$ scaling regime. We thus excluded these results in determining the continuum estimate of $\chi_{\mathrm{disc}}$, thereby using only the $N_{\tau}=12,16$ results. The fact that this difference in the scaling behaviour shows up in the magnitude of the disconnected chiral susceptibility and is not reflected in the location of its peak is not unfamiliar and shows up also in other fluctuations observables, e.g. cumulants of conserved charge fluctuations. Here it can be addressed to large cut-off effects arising from taste violations inherent to staggered fermion actions. This may also be the case for the disconnected chiral susceptibility as it is known that in particular at low temperatures the chiral condensate is dominated by contributions from the pseudo-scale meson sector [60], where taste violation effects are large. Furthermore, the data for $m_{s}^{2}\,\chi_{\mathrm{disc}}/T_{c}^{4}$ for different lattice cutoffs obtained using HISQ discretization thus shows that taste violation effects strongly influence chiral symmetry breaking at finite lattice spacing, resulting in the presence of large cutoff effects in observables sensitive to the pseudo-scalar meson sector of QCD. The comparison between different fermion discretizations also highlights the importance of calculating chiral observables using lattice fermion discretizations, e.g. MDWF, where the chiral symmetry breaking and finite cutoff effects are largely disentangled.
@@
Furthermore, our MDWF results for $\chi_{\mathrm{disc}}$ lie within the continuum estimate band obtained using the HISQ data, even though our calculations with MDWF were performed on a finite cutoff i.e. an $N_{\tau}=8$ lattice. Of course, results for physical observables obtained by using different fermion discretizations should agree in the continuum limit. The fact that our results obtained using MDWF on a finite lattice spacing already agree with the continuum estimate of the HISQ results indicates that finite cutoff effects are subdominant for finer lattices, i.e. for $N_{\tau}>8$. While it would require a lattice study for explicit verification of this observation, from Eq. (9) it is clear that the additional term contained in the mass derivative of the MDWF operator becomes irrelevant on finer lattices, as its magnitude decreases with decreasing the lattice spacings.
@@
Next, we compare how well the topological tunnelings that occur in QCD are represented in the gauge configurations generated using the HISQ and MDWF discretizations for fermions. This comparison is important for two reasons. First, for the domain-wall fermions, one can be assured that the topological tunnelings are adequately represented even at non-vanishing lattice spacing without any significant contributions arising from localized dislocations whose size is of the order of a single lattice spacing, quite unlike in the case for staggered fermions at a similar value of the lattice cutoff. Secondly, the domain-wall Dirac matrix violates continuum chiral symmetries through terms which are irrelevant in the continuum limit, unlike the HISQ discretization where the chiral symmetry group and hence the notion of an index is altered on a finite lattice due to rooting. It is not apriori clear what are the consequences of this fact for the topological susceptibility obtained using the HISQ discretization, i.e. how fine lattices would be required to perform a correct continuum extrapolation of this observable. In the literature there have been extensive discussions on whether the topological fluctuations are correctly accounted for in rooted staggered fermions [61]. In the weak-coupling lattice perturbation theory it has been shown that a four flavor-taste staggered determinant factors into four equivalent determinants and that the correct flavor symmetries emerge in the continuum limit [62, 63, 64, 65]. For staggered fermions, however, the flavor singlet part of chiral symmetry which is sensitive to topology, has to be defined carefully on a finite lattice. Previous studies reconstructing the t’Hooft vertex and the flavor singlet meson correlator from the near-zero-mode quartets, showed the emergence of the expected behavior of these observables in the continuum [66] and more recently reproduced the $\eta^{\prime}$ meson mass using the staggered discretization of QCD in the continuum limit [67].

### v2

@@ IV Results
IV.1 Estimating the pseudocritical temperature
@@
for which we present our results in Fig. 1. Since the transition associated with the restoration of the nonsinglet $\mathrm{SU}_{A}(2)$ symmetry is a smooth crossover in QCD with physical quark masses, this observable has a rather smooth dependence on the temperature. The chiral susceptibility, on the other hand, is known to be more sensitive to the restoration of symmetries and, in particular, to diverge in the chiral limit. This also holds true for its disconnected part, whereas its connected part is expected to diverge only if the $\mathrm{U}_{A}(1)$ symmetry is restored.
FIG. 1: The subtracted chiral condensate $\Delta_{ls}$ normalized by $(T/T_{pc})^{4}$ is shown as a function of the temperature $T$ calculated using MDW fermions (red pentagons) on an $N_{\tau}=8$ lattice and compared with results from Ref. [58] obtained using the HISQ discretization on $N_{\tau}=8,12,16$ lattices. For the normalization factor $T_{pc}$ we used the pseudocritical temperature $T_{pc}=155\,\mathrm{MeV}$ reported in the earlier work [9].
The results for the disconnected part of the chiral susceptibility as a function of the temperature $T$ is shown in Fig. 2 (top). This observable shows a pronounced peak in a small range of temperatures and falling off both at lower and higher temperatures. The location of this peak defines a pseudocritical temperature, which we denote as $T_{pc,N_{\tau}=8}$.
For the determination of the pseudocritical temperature, we normalized the data points of the disconnected part of the chiral susceptibility in the temperature-independent form $m_{s}^{2}\,\chi_{\mathrm{disc}}/T_{pc}^{4}$, using the temperature values determined using the $r_{0}$-scale and the pseudocritical temperature $T_{pc}$ from Refs. [9, 55], which are also given in Table 1. The normalized data points are interpolated using a Padé fit of order $[1/3]$. From the zero of the derivative of this rational function with respect to the temperature, shown in Fig. 2 (bottom), we have estimated the pseudocritical temperature to be
@@
which we henceforth refer to as $T_{pc}$. This is consistent with the earlier estimate of the pseudocritical temperature, $155(9)\,\mathrm{MeV}$, given in Ref. [9], with the improvement of reduced statistical errors. Our estimate of $T_{pc}$ is also consistent with the corresponding ones obtained using the HISQ and stout actions after performing a continuum extrapolation, which are $156.5(1.5)\,\mathrm{MeV}$ [4] and $158.0(6)\,\mathrm{MeV}$ [6], respectively. Furthermore, $\chi_{\text{disc}}$ becomes small only at temperatures $T>186\,\mathrm{MeV}$, signalling a significant contribution from the topological fluctuations of the QCD vacuum in the vicinity of $T_{pc}$ which we discuss in detail in the next section.
FIG. 2: The disconnected part of the chiral susceptibility shown as a function of $T$ (top) and the extraction of the pseudocritical temperature $T_{pc,N_{\tau}=8}$ (black point) from the point where the slope of $m_{s}^{2}\chi_{\text{disc}}$ goes to zero (bottom). For the normalization factor $T_{pc}$ we used the pseudocritical temperature $T_{pc}=155\,\mathrm{MeV}$ reported in the earlier work [9].
@@
Results for this mixed susceptibility are shown in Fig. 3. Our data, obtained with the MDWF action on $N_{\tau}=8$ lattices, are compared with corresponding results obtained in calculations with the HISQ action on lattices with identical spatial and temporal lattice extent [59]. We note that results obtained within the MDWF and HISQ discretization schemes are in good quantitative agreement. Although the statistical noise on this observable is relatively large in the MDWF calculation, it also shows a peak at $T^{\chi_{us}}_{N_{\tau}=8}\simeq 160\,\mathrm{MeV}$, which is in good agreement with the quoted HISQ result $T^{\chi_{us}}_{\mathrm{HISQ}}=159.39(22)\,\mathrm{MeV}$ [59]. Also the magnitudes agree well despite the different discretization errors arising in the MDWF and HISQ discretization schemes, respectively. As discussed in more detail in Sec. IV.3, this is quite different from the large difference observed in calculations of the disconnected chiral susceptibility on lattices with identical temporal extent, $N_{\tau}=8$. In the context of calculations with staggered fermions, this is understood to arise from taste symmetry violations which are much more important in the light quark sector than in the strange quark sector. This also reaffirms the observation made in [59] that the strange quark mass is relatively heavy and does not break the almost exact two-flavor chiral symmetry in the QCD action due to the light quarks explicitly. It therefore appears as an external parameter in the energylike scaling variable used in the scaling analysis in the vicinity of the chiral transition, just like the chemical potentials corresponding to the conserved charges.
@@
FIG. 3: The mixed-mass susceptibility $\chi_{us}$ as a function of the temperature $T$. Shown are results from calculations with the MDWF action (red pentagons) and the HISQ action (blue squares) (taken from Ref. [59]) on lattices with temporal extent $N_{\tau}=8$. The band is obtained from a Padé fit to the MDWF data. We have normalized the data with a factor of $T_{pc}^{4}$, using $T_{pc}=158.7\,\mathrm{MeV}$ as given in Eq. (14).
@@
Having discussed the restoration of the nonsinglet part of chiral symmetry, it would be also interesting to check how well the chiral Ward identities are followed by the MDWF action, even at finite lattice spacing.
IV.2 The fate of the singlet and nonsinglet parts of chiral symmetry
@@
Once the nonsinglet $\mathrm{SU}_{A}(2)$ part of chiral symmetry is restored at $T_{pc}$, one would then ask the question, when the singlet $\mathrm{U}_{A}(1)$ subgroup of the chiral symmetry is effectively restored. We now want to estimate the magnitude of the $\mathrm{U}_{A}(1)$ breaking at temperatures greater than $T_{pc}$, which is done through a calculation of the observable $\chi_{\pi}-\chi_{\delta}$ as motivated earlier. According to the chiral Ward identities, the integrated pion correlator $\chi_{\pi}$ should be related to the chiral condensate as seen in Eq. (7). On the other hand, the connected piece of the chiral susceptibility is just the scalar isotriplet correlator, $\chi_{\text{conn}}\equiv\chi_{\delta}$. Previous studies [24, 48] with domain-wall fermions have reported $\chi_{\pi}-\chi_{\delta}$ to be finite at $196\,\mathrm{MeV}$, anticipating that the $\mathrm{U}_{A}(1)$ symmetry will be effectively restored beyond this temperature [9]. A different study, also with another variant of domain-wall fermions, but on relatively small lattice volumes, reports on the effective restoration of the $\mathrm{U}_{A}(1)$ symmetry at temperatures around $T_{c}$ [60].
@@
In order to address this apparently unsettled issue, we calculate the temperature dependence of the topological susceptibility to have an independent measure of the $\mathrm{U}_{A}(1)$ breaking. We recall that the restoration of the nonsinglet part of chiral symmetry can also be inferred from Eq. (12). Moreover, the finite volume correction to $\chi_{\mathrm{top}}$ is dominated by the $\eta^{\prime}$ meson mass rather than the pion, and is expected to be mild at temperatures beyond $T_{pc}$. We obtained $\chi_{\text{top}}$ on the gauge configurations generated with MDWF, by calculating the average variance of the number of zero modes counted using as probe the overlap Dirac operator, details of which are mentioned in Sec. III. Our results for $\chi_{\text{top}}$ as a function of the temperature are shown in Fig. 4. From next-to-leading order (NLO) chiral perturbation theory, the topological susceptibility can be derived to be
$\chi_{\text{chiPT}}^{\text{NLO}}$ $=\frac{z}{(1+z)^{2}}M^{2}_{\pi}F^{2}_{\pi}[1+\delta_{1}],\;$ (16)
@@
where $\delta_{1}=\frac{2M_{\pi}}{F^{2}_{\pi}}(h_{1}^{r}-h_{3}^{r}-l_{4}^{r}-l_{7})$ in terms of low energy constants of SU(2) chiral perturbation theory [61]. For our case, $z$, which is the ratio of mass of up and down quarks, is simply $z=1$. Using the latest values of $F_{\pi}=94.15\,\mathrm{MeV}$, $M_{\pi}=134.98\,\mathrm{MeV}$ and the quantity $(h_{1}^{r}-h_{3}^{r}-l_{4}^{r}-l_{7})=-0.0114$ from the FLAG review [62], we obtain the $(\chi_{\text{chiPT}}^{\text{NLO}})^{1/4}=78.8\pm 2.9\,\mathrm{MeV}$, where the error predominantly comes from the error in the determination of $F_{\pi}$. As discussed in the previous section, the contribution to the chiral observables from strange quarks is small compared to the light quarks. Hence, we do not consider three-flavor chiral perturbation theory. Our results for the topological susceptibility in (2+1)-flavor QCD (triangles) with MDWF start to deviate from the NLO chiral perturbation theory prediction shown as a band in Fig. 4 below $T_{pc}$, alluding toward an effective restoration of the nonsinglet part of chiral symmetry. The deviation from the chiral perturbation theory result becomes larger with increasing temperatures, showing a $\approx 50\%$ deviation at the highest temperature of $186\,\mathrm{MeV}$. However the magnitude of $\chi_{\text{top}}$ at temperatures higher than $T_{pc}$ successively gets closer to the values of $m_{\ell}^{2}\,\chi_{\text{disc}}$, besides a small deviation resulting from the finite mass of the light quarks used in this work, signalling an effective restoration of the nonsinglet part of chiral symmetry.
@@
Furthermore, the temperature dependence of our data for $\chi_{\text{top}}^{1/4}/T_{pc}$ at $T>160\,\mathrm{MeV}$ can be characterized by $(T/T_{pc})^{-\alpha}$ with $\alpha=2.8(1)$. The exponent $\alpha\approx 2$ describes the temperature dependence of the topological susceptibility in QCD with two light quark flavors within a dilute instanton gas model, which assumes fluctuations of isolated uncorrelated instantons, and should persist at asymptotically high temperatures. The fact that the exponent $\alpha>2$, reflects a nontrivial breaking of the $\mathrm{U}_{A}(1)$ symmetry up to $T\approx 186\,\mathrm{MeV}$, which is unlikely due to a dilute instanton gaslike scenario. It is also known from a recent study with the HISQ discretization that the $\mathrm{U}_{A}(1)$ symmetry is effectively restored at $T\gtrsim 180\,\mathrm{MeV}$, after performing a continuum extrapolation [34]. This is consistent with the fact that we do not observe an effective restoration of the $\mathrm{U}_{A}(1)$ part of chiral symmetry up to the highest temperature we have studied using MDWF, which is $T=186\,\mathrm{MeV}$.
IV.3 Comparing chiral observables in the light quark sector calculated with MDWF and HISQ
@@
Having obtained precise results on the chiral observables using MDWF, even on a finite $N_{\tau}=8$ lattice, it is imperative to determine how well these results compare with results obtained using different fermion discretizations. The staggered fermion discretization scheme explicitly breaks the nonsinglet flavor symmetries on the lattice and only respects a $\mathrm{U}(1)$ subgroup of the whole chiral symmetry group of QCD with two light quark flavors, which is an admixture of the physical flavor and the taste symmetries. It is known that the chiral symmetries in terms of the chiral Ward identities are recovered in the continuum limit for observables calculated using the HISQ discretization. This has been shown, for example, through the exact matching of the disconnected part of the chiral susceptibility and the topological susceptibility in the continuum limit for physical gauge ensembles in (2+1)-flavor QCD generated with the HISQ action [29, 34]. On the other hand, in earlier work with MDWF in Ref. [9], it was observed that $\chi_{\text{disc}}$ calculated using domain-wall fermions was $50\%$ larger than the result computed using the staggered discretization in the crossover region. It would thus be important to now address this puzzle, especially since we know that for a meaningful comparison one has to perform a continuum extrapolation of the staggered results.
@@
FIG. 5: $m_{s}^{2}\,\chi_{\mathrm{disc}}/T_{pc}^{4}$ computed with the MDWF action for $N_{\tau}=8$, in comparison to results obtained using the HISQ action. The continuum estimate was computed from the HISQ $N_{\tau}=12,16$ results only. For the normalization factor $T_{pc}$ we used the pseudocritical temperature $T_{pc}\equiv T_{pc,N_{\tau}=8}$ given in Eq. (14).
@@
Results for the observable $\chi_{\text{disc}}$ as a function of the temperature are shown in Fig. 5, comparing the results obtained using the HISQ discretization for lattice spacings corresponding to $N_{\tau}=8,12,16$ with our results calculated using MDWF on an $N_{\tau}=8$ lattice. From a continuum extrapolation of the peak positions associated with the HISQ $N_{\tau}=8,12,16$ data for $\chi_{\mathrm{disc}}$, we have estimated the pseudocritical temperature with an $\mathcal{O}(1/N_{\tau}^{2})$ fit resulting in $T_{pc}^{\mathrm{HISQ}}=157.4(7)\,\mathrm{MeV}$, which agrees well within errors with our corresponding MDWF $N_{\tau}=8$ estimate of $T_{pc,N_{\tau}=8}=158.7{}_{{}\mathop{-}2.3}^{{}\mathop{+}2.6}\,\mathrm{MeV}$, as given in Eq. (14). Unlike the location of its peaks, the numerical values of $\chi_{\mathrm{disc}}$ obtained using HISQ action on an $N_{\tau}=8$ lattice are clearly not within the $\mathcal{O}(a^{2})$ scaling regime. We thus excluded these results in determining the continuum estimate of $\chi_{\mathrm{disc}}$, thereby using only the $N_{\tau}=12,16$ results. The fact that this difference in the scaling behavior shows up in the magnitude of the disconnected chiral susceptibility and is not reflected in the location of its peak is not unfamiliar and shows up also in other fluctuations observables, e.g. cumulants of conserved charge fluctuations. Here it can be addressed to large cutoff effects arising from taste violations inherent to staggered fermion actions. This may also be the case for the disconnected chiral susceptibility as it is known that in particular at low temperatures the chiral condensate is dominated by contributions from the pseudo-scalar meson sector [63], where taste violation effects are large. Furthermore, the data for $m_{s}^{2}\,\chi_{\mathrm{disc}}/T_{c}^{4}$ for different lattice cutoffs obtained using HISQ discretization thus shows that taste violation effects strongly influence chiral symmetry breaking at finite lattice spacing, resulting in the presence of large cutoff effects in observables sensitive to the pseudo-scalar meson sector of QCD. The comparison between different fermion discretizations also highlights the importance of calculating chiral observables using lattice fermion discretizations, e.g. MDWF, where the chiral symmetry breaking and finite cutoff effects are largely disentangled.
@@
Furthermore, our MDWF results for $\chi_{\mathrm{disc}}$ lie within the continuum estimate band obtained using the HISQ data, even though our calculations with MDWF were performed on a finite cutoff i.e., an $N_{\tau}=8$ lattice. Of course, results for physical observables obtained by using different fermion discretizations should agree in the continuum limit. The fact that our results obtained using MDWF on a finite lattice spacing already agree with the continuum estimate of the HISQ results suggests that finite cutoff effects are subdominant for finer lattices, i.e., for $N_{\tau}>8$. In fact, from Eq. (9) it is evident that the additional term contained in the mass derivative of the MDWF operator becomes irrelevant on finer lattices, as its magnitude decreases with decreasing the lattice spacings. Nonetheless, in order to finally arrive at a well controlled comparison between different discretization schemes this should eventually be verified explicitly in a direct lattice QCD calculation within the MDWF discretization scheme using larger temporal lattice extents, $N_{\tau}$.
@@
Next, we compare how well the topological tunnelings that occur in QCD are represented in the gauge configurations generated using the HISQ and MDWF discretizations for fermions. This comparison is important for two reasons. First, for the domain-wall fermions, one can be assured that the topological tunnelings are adequately represented even at non-vanishing lattice spacing without any significant contributions arising from localized dislocations whose size is of the order of a single lattice spacing, quite unlike in the case for staggered fermions at a similar value of the lattice cutoff. Second, the domain-wall Dirac matrix violates continuum chiral symmetries through terms which are irrelevant in the continuum limit, unlike the HISQ discretization where the chiral symmetry group and hence the notion of an index is altered on a finite lattice due to rooting. It is not apriori clear what are the consequences of this fact for the topological susceptibility obtained using the HISQ discretization, i.e., how fine lattices would be required to perform a correct continuum extrapolation of this observable. In the literature there have been extensive discussions on whether the topological fluctuations are correctly accounted for in rooted staggered fermions [64]. In the weak-coupling lattice perturbation theory it has been shown that a four flavor-taste staggered determinant factors into four equivalent determinants and that the correct flavor symmetries emerge in the continuum limit [65, 66, 67, 68]. For staggered fermions, however, the flavor singlet part of chiral symmetry which is sensitive to topology, has to be defined carefully on a finite lattice. Previous studies reconstructing the t’Hooft vertex and the flavor singlet meson correlator from the near-zero-mode quartets, showed the emergence of the expected behavior of these observables in the continuum [69] and more recently reproduced the $\eta^{\prime}$ meson mass using the staggered discretization of QCD in the continuum limit [70].